Determine the tangent to the curve 3y2 = x3 at (3,3) and calculate the area of the triangle bounded by the tangent line, the x-axis and the line x=3.


1
Expert's answer
2021-06-14T18:18:30-0400

Let us determine the tangent to the curve "3y^2 = x^3" at "(3,3)". Since "6yy'=3x^2," we conclude that "y'=\frac{x^2}{2y}," and hence "y'(3)=\frac{3^2}{2\cdot 3}=\frac{3}{2}." It follows that the equation of the tangent to the curve is "y-3=\frac{3}{2}(x-3)," which is equivalent to "y=\frac{3}{2}x-\frac{3}{2}."


Let us calculate the area of the triangle bounded by the tangent line, the x-axis and the line "x=3". If "y=0," then "\frac{3}{2}x-\frac{3}{2}=0," and hence "x=1." If "x=3," then "y=\frac{3}{2}\cdot 3-\frac{3}{2}=3." We conclude that the vertices of a triangle are "A(1,0),\ B(3,0)" and "C(3,3)." It follows thta this triangle is right with legs "3-1=2" and "3". Therefore, its area is "\frac{1}{2}\cdot2\cdot3=3."